OpenStudy (anonymous):

The soccer team collected $800 at a car wash fundraiser. They charged $5.00 for small vehicles and $10.00 for larger vehicles. The amount collected can be modeled by the equation 5x + 10y = 800, where x represents the number of small vehicles and y represents the number of larger vehicles. If the number of larger vehicles washed was 50, how many small vehicles were washed in total?

OpenStudy (anonymous):

60 small vehicles because 10 * 50 =500 800-500=300 300/5=60 small cars